Factor..show steps tjank u
12xy+8x-15y-10

Answer by tommyt3rd(5050)

12xy+8x-15y-10=
(12xy+8x)-(15y+10)=
4x(3y+2)-5(3y+2)=
(4x-5)(3y+2)


:)